OpenStudy (anonymous):

Select the ordered pair from the choices below that is a solution to the following system of equations: 4y = 2x + 10 8x − 3y = -14 a. (-1, 2) b. (7, 6) c. (-3, 2) d. (5, 5)

OpenStudy (anonymous):

a. (-1, 2)

OpenStudy (anonymous):

@El-naggar : how ?

OpenStudy (anonymous):

sorry , i'll tell you how to solve these two equations

OpenStudy (anonymous):

multiply the first equation by 4 it gives 16y=8x+40 rearranging the equations 8x=16y-40 , 8x=3y-14 subtracting the two equations gives 0=13y-26 then y=2 hence x can be calculated from the first equation x= ((4*2)-10)/2 = -1